OpenStudy (anonymous):

Solve the following system of equations. 2x=6 3x-y=9 and my answer are (0,3) (-3,0) (0,0) (3,0)

OpenStudy (judygreeneyes):

2X =6 3X -Y=9 Multiply 1st equation by 3 on both sides, and 2nd equation by -2 on both sides (this gives us equations that are easier to solve. 6X = 18 -6X+2Y = -18 Now add the equations together: 6X - 6X = 0, so they cancel out: New equation: 2Y = 0 Y = 0 Now plug Y=0 into the second equation: 3X-Y=9 3X - 0 = 9 3X=9 X = 3 The (x,y) answer is (3,0)